# ResQuNNs:Towards Enabling Deep Learning in Quantum Convolution Neural Networks

### v4

@@ V-B Three Quanvolutional Layers
VI Conclusion
In this paper, we presented a novel framework for developing Quanvolutional Neural Networks (QuNNs) by introducing trainable quanvolutional layers and the concept of Residual Quanvolutional Neural Networks (ResQuNNs) for enabling the deep learning in QuNNs. Our approach addresses the significant challenges of gradient flow and adaptability in deep QuNN architectures, enabling enhanced training performance and scalability.
A key contribution of this work is the development of trainable quanvolutional layers, which enhance the learning potential of QuNNs, and the strategic incorporation of residual blocks to ensure comprehensive gradient accessibility across the network in multi-layered QuNNs, thus, enabling deeper, more complex QuNNs architectures to be effectively trained. Our empirical findings highlight the importance of optimal placement of residual blocks, providing critical insights for future QuNN architecture designs with even more deeper layers.
In summary, our work signifies a substantial leap in quantum deep learning, offering new directions for theoretical and practical quantum computing research. By tackling core issues of adaptability and gradient flow, our work paves the way for the development of more robust and efficient quantum neural networks, marking a pivotal step toward harnessing the full potential of quantum technologies.

### v5

@@ 5.2 Three Quanvolutional Layers
6 Conclusion
In this paper, we presented a novel framework for developing Quanvolutional Neural Networks (QuNNs) by introducing trainable quanvolutional layers and the concept of Residual Quanvolutional Neural Networks (ResQuNNs) for enabling the deep learning in QuNNs. Our approach addresses the significant challenges of gradient flow and adaptability in deep QuNN architectures, enabling enhanced training performance and scalability. A key contribution of this work is the development of trainable quanvolutional layers, which enhance the learning potential of QuNNs, and the strategic incorporation of residual blocks to ensure comprehensive gradient accessibility across the network in multi-layered QuNNs, thus, enabling deeper, more complex QuNNs architectures to be effectively trained. Our empirical findings highlight the importance of optimal placement of residual blocks, providing critical insights for future QuNN architecture designs with even more deeper layers. In summary, our work signifies a substantial leap in quantum deep learning, offering new directions for theoretical and practical quantum computing research. By tackling core issues of adaptability and gradient flow, our work paves the way for the development of more robust and efficient quantum neural networks, marking a pivotal step toward harnessing the full potential of quantum technologies.